OpenStudy (anonymous):

Stobart puts $40 in a bank where the interest rate is 6% a year compounded continuously. a) What will be the value of the investment in 12 years? b)How long will it take for the investment to grow to $100?

OpenStudy (gschibby):

Value year 1 = 40 * 1,06 Value year 2 = 40 * 1,06^2 Value year n = 40 * 1,06^n Value year 12 = ? ;)

OpenStudy (anonymous):

Value year 40*e^(0.06) 12 ?

OpenStudy (gschibby):

Value year 12 = 40 * 1,06^12 ;)

OpenStudy (gschibby):

80,5

OpenStudy (gschibby):

For the other part you plot in 100 = 40 * 1,06^x and solve for x

OpenStudy (anonymous):

why do you add 1,06 isn't it 0.06

OpenStudy (gschibby):

No, 0.06 is just the increase. I guess you could type it like this: 40+(40*0.06) But it's much easier this way :)

OpenStudy (gschibby):

And it's not + it's *, you multiply with 1.06 ;)

OpenStudy (gschibby):

\[ValueYear12 = 40 * 1.06^{12}\]